OpenStudy (anonymous):

Differentiate: G(x)=sqrt(1-x^2) * arccos(x)

OpenStudy (anonymous):

I got \[G'(x)=-\frac{ \arccos(x) }{ \sqrt{1-x^2} }-1\] but the book says there should be an extra x in the numerator

hartnn (hartnn):

know about the chain rule ?

OpenStudy (anonymous):

Yes, I'm guessing that's where I slipped

hartnn (hartnn):

d/dx (sqrt(1-x^2)) = 1/2 sqrt (1-x^2) times d/dx (1-x^2) and whats d/dx (1-x^2) = ...?

OpenStudy (anonymous):

I thought d/dx(1-x^2) is just 2 no?

OpenStudy (anonymous):

I mean -2

hartnn (hartnn):

d/dx x^2 = ... ?

OpenStudy (anonymous):

*facepalm* 2x

OpenStudy (anonymous):

I was thinking the x turns into a 1 and leaves 2 alone, thanks

hartnn (hartnn):

welcome ^_^